OpenStudy (anonymous):

Salim sells fax machines that come in Standard and Enhanced models. The Standard model sells for $325 and the Enhanced model sells for $385. If Salim sold a total of 35 units and took in $12,575, how many of the Enhanced model did he sell?

OpenStudy (anonymous):

S + E = 35 325 S + 385 E = 12,575 Either substitute E = (35-S) into the second equation or multiply the first by -325 and add the two and you are nearly done.